The length of a rectangle is 6 cm more than 1 1/2 times the width. The perimiter of the rectangle is 56 cm. Find the dimensions of the rectangle.Can someone show me the work too?
Answer by stanbon(75887)

The length of a rectangle is 6 cm more than 1 1/2 times the width. The perimiter of the rectangle is 56 cm. Find the dimensions of the rectangle
-------------------------------------------------
Perimeter = 2(W + L)
56 = 2(W + (3/2)W + 6)
28 = (5/2)W + 6
(5/2)W = 22
W = 44/5 = 8.8 cm
--------------------------
L = (3/2)W + 6
L = (3/2)*8.8 + 6
L = 19.2 cm
===================
Cheers,
Stan H.